OpenStudy (anonymous):

Sugar is put into a large quantity of water and the mixture is stirred. After 2 minutes, 50% of the sugar has dissolved. How much longer will it take until 90% of the sugar has dissolved?

OpenStudy (anonymous):

Answer is 4.644 mins longer (eg 6.644 minutes total). How to solve?

OpenStudy (misty1212):

and more to the point, how could we know the answer? you got a formula or sommat?

OpenStudy (anonymous):

A(t) = A0xe^kt? I'm not sure, my teacher didn't go over this topic that well, so I'm extremely clueless.

OpenStudy (mathstudent55):

Every 2 minutes, half of the sugar dissolves. Start with 100% undissolved sugar at t = 0 At t = 2 min, 50% is still undissolved. At t = 2, 0.5 is undissolved At t = 4 min, 25% is still undissolved. At t = 4, 0.5 * 0.5 = 0.25 is undissolved At t = 6 min, 12.5% is still undissolved. At t = 6, 0.5 * 0.5 * 0.5 = 0.125 is undissolved At t = x 2-minute periods, 0.5^x is undissolved 0.1 = 0.5^x log 0.1 = x log 0.5 x = log 0.1/log 0.5 x = 3.3219 Since x is the number of 2-minute periods, the time is 2 * 3.3219 min = 6.6438 minutes

OpenStudy (mathstudent55):

6.6438 minutes is the total time for 90% of the sugar to dissolve. The question is how much more time than the 2 minutes will it take, so the answer is 4.638 minutes.

OpenStudy (anonymous):

Doesn't the info say that 50% has dissolved? Does it imply that it decreases by half every time?

OpenStudy (anonymous):

Yeah, that's what I was thinking.

OpenStudy (anonymous):

But wouldn't it completely dissolve by 4 minutes, then? Since it decreased by half the first time. 0.0

OpenStudy (mathstudent55):

If 50% dissolves each 2 minutes, then 50% of the sugar remains undissolved each period.

OpenStudy (mathstudent55):

No. Each 2 minutes that pass, half of the remaining undissolved sugar dissolves.

OpenStudy (anonymous):

Okay. I see.